# PDE-Transformer: A Continuous Dynamical Systems Approach to Sequence Modeling

### v2

@@ B.3 Supplementary Analysis
B.3.2 Robustness and Sensitivity Analysis.
@@
B.3.3 Frequency Domain Analysis.
@@
B.4 Core Mechanism and Theoretical Analysis
B.4.1 Theoretical Design Principles.
This seemingly simple formula embodies two key, principled design choices:
  1. 1.
    Update Form Based on Numerical Analysis. The residual form $X+\dots$ is not merely a conventional skip connection. In numerical analysis, it can be rigorously interpreted as a single-step forward Euler method for solving the diffusion equation $\frac{\partial u}{\partial t}=\alpha_{\text{coeff}}\nabla^{2}u$, where the time step $\Delta t$ is effectively absorbed into the learnable coefficient $\alpha$. According to Theorem 1.2 (Exponential Decay of the Gradient Norm), this update form guarantees that the system’s energy gradient converges at an exponential rate, providing a solid mathematical foundation for stability:
    $\left\|\frac{\delta E}{\delta u}(t+\Delta t)\right\|_{L^{2}}\leq\left\|\frac{\delta E}{\delta u}(t)\right\|_{L^{2}}e^{-\mu\alpha}$ (10)
    where $\mu$ is a positive definite constant of the system, ensuring the stability of the information propagation process.
  2. 2.
    Learnable Adaptive Diffusion Strength. The diffusion coefficient $\alpha$ is a learnable scalar parameter, which endows the model with a crucial adaptive capability. Based on Theorem 3.1 (Frequency Properties of Multi-Scale Filters), the learning process of $\alpha$ is, in fact, an optimization of the layer’s frequency response function:
    $\hat{\Delta}_{\text{adaptive}}(\omega)=\alpha\cdot\left(-\frac{4}{h^{2}}\sin^{2}\left(\frac{\omega h}{2}\right)\right)$ (11)
    The model can autonomously decide the required intensity of information smoothing at each layer based on the task and data, and can even effectively ”turn off” the layer by learning an $\alpha$ close to zero.
B.4.2 Theoretical Complementarity Analysis.
The PDE Diffusion Layer introduces a powerful inductive bias of local smoothness into the Transformer model. It forms a profound theoretical complementarity with the self-attention mechanism:
  • Self-Attention: Performs sparse, content-based global information retrieval.
  • PDE Diffusion: Performs dense, structure-based local information integration.
According to Theorem 2.1 (Polynomial Decay of the Heat Kernel), this complementarity is manifested in a fundamental difference in their information propagation patterns. For two positions at a distance of |xy||x-y|, their influence decay patterns are:
  • Attention Weights (Typical): $A(x,y)\sim e^{-c|x-y|}$ (Exponential decay)
  • Diffusion Kernel: $K_{t}(x,y)\sim|x-y|^{-1}$ (Polynomial decay)
Polynomial decay provides a stronger long-range connection capability, effectively supplementing the deficiencies of standard attention mechanisms in capturing ultra-long-range dependencies.
B.4.3 Inherent Positional Awareness.
The PDE Diffusion Layer provides the model with an inherent, structural sense of position that is orthogonal to traditional positional encodings. While standard Transformers rely on externally injected signals (e.g., sinusoidal encodings) to perceive order, our diffusion mechanism, through its intrinsic local stencil, tightly couples the representation of each token to its immediate neighbors, thus naturally encoding the topology and relative order of the sequence.

### v3

@@ A.3 Supplementary Analysis
A.3.2 Robustness and Sensitivity Analysis.
@@
A.3.3 Frequency Domain Analysis.
@@
Appendix B Detailed Theoretical Proofs
This appendix provides self-contained proofs of the main stability statements for the PDE diffusion operator used in the paper. Throughout, we use Neumann (zero-flux) boundary conditions implemented by
emphreplication at the endpoints (often called replicate or nearest padding in deep-learning libraries):
$u_{0}=u_{1},\qquad u_{L+1}=u_{L}.$ (17)
This convention is consistent with the discrete Neumann Laplacian defined below. The reflect padding rule corresponds to a different ghost-point convention and would change the boundary rows; throughout, our analysis and implementation assume Eq. (17).
B.1 Notation and the discrete Neumann Laplacian
Let $L$ be the sequence length and $d$ the hidden dimension. Representations are matrices $X\in\mathbb{R}^{L\times d}$, with channels indexed by $j\in\{1,\dots,d\}$. The 1D discrete Neumann Laplacian $\Delta_{N}\in\mathbb{R}^{L\times L}$ acts on a vector $x\in\mathbb{R}^{L}$ by
$(\Delta_{N}x)_{i}=\begin{cases}x_{2}-x_{1},&i=1,\\ x_{i-1}-2x_{i}+x_{i+1},&2\leq i\leq L-1,\\ x_{L-1}-x_{L},&i=L.\end{cases}$ (18)
For matrix inputs $X$, $\Delta_{N}$ acts channel-wise along the sequence dimension: $(\Delta_{N}X)_{\cdot j}=\Delta_{N}(X_{\cdot j})$.
We define the diffusion operator
$\mathcal{S}_{\alpha}(X)\triangleq(I+\alpha\Delta_{N})X,\qquad 0\leq\alpha<\tfrac{1}{2}.$ (19)
Unless stated otherwise, $\|\cdot\|_{2}$ denotes the spectral norm and $\|\cdot\|_{F}$ denotes the Frobenius norm.
B.2 Proof of the spectrum of $\Delta_{N}$
Lemma B.1 (Spectrum of the discrete Neumann Laplacian).
$\Delta_{N}$ is real symmetric and negative semi-definite. Its eigenvalues are
$\lambda_{k}=-4\sin^{2}\!\Bigl(\frac{\pi k}{2L}\Bigr),\qquad k=0,1,\dots,L-1,$ (20)
so that $\lambda_{k}\in[-4,0]$ and $\lambda_{0}=0$ corresponds to the constant mode.
Proof.
(1) Symmetry and semi-definiteness. From Eq. (18), $\Delta_{N}$ is tri-diagonal with boundary rows $(\Delta_{N}x)_{1}=x_{2}-x_{1}$ and $(\Delta_{N}x)_{L}=x_{L-1}-x_{L}$, hence its matrix form is symmetric. Moreover, for any $x\in\mathbb{R}^{L}$,
[…]
which implies $\Delta_{N}\preceq 0$.
(2) Eigenpairs. Consider the cosine family (a DCT-type basis)
$v^{(k)}_{i}=\cos\!\left(\frac{\pi k}{L}\left(i-\frac{1}{2}\right)\right),\qquad i=1,\dots,L,$ (22)
for $k=0,1,\dots,L-1$. A direct computation for interior indices uses the identity $\cos(a-b)-2\cos(a)+\cos(a+b)=2(\cos b-1)\cos a$ with $b=\tfrac{\pi k}{L}$ to give
$(\Delta_{N}v^{(k)})_{i}=2\Bigl(\cos\!\bigl(\tfrac{\pi k}{L}\bigr)-1\Bigr)v^{(k)}_{i}=-4\sin^{2}\!\Bigl(\tfrac{\pi k}{2L}\Bigr)v^{(k)}_{i},\qquad 2\leq i\leq L-1.$ (23)
The same eigenvalue also satisfies the boundary rows in Eq. (18) (consistent with the endpoint replication convention), hence each $v^{(k)}$ is an eigenvector with eigenvalue $\lambda_{k}$ in Eq. (20).
(3) Range. For $k=0$, $\lambda_{0}=0$ and $v^{(0)}$ is constant. For $k\geq 1$, $\sin(\cdot)>0$ implies $\lambda_{k}<0$, and the minimum approaches 4-4. This completes the proof. ∎
B.3 Proof of spectral non-expansiveness and strict contraction off the nullspace
Theorem B.2 (Spectral non-expansiveness; strict contraction off the nullspace).
For $0\leq\alpha<\tfrac{1}{2}$, the diffusion operator $\mathcal{S}_{\alpha}$ in Eq. (19) satisfies
$\|\mathcal{S}_{\alpha}\|_{2}\leq 1,\qquad\|\mathcal{S}_{\alpha}(X)\|_{F}\leq\|X\|_{F}\ \ \text{for all }X\in\mathbb{R}^{L\times d}.$ (24)
Let $\Pi_{\perp}$ be the orthogonal projector onto $\mathrm{Null}(\Delta_{N})^{\perp}$ (i.e., the zero-mean subspace along the sequence dimension). Then there exists $\kappa(\alpha,L)\in(0,1)$ such that
$\|\Pi_{\perp}\mathcal{S}_{\alpha}(X)\|_{F}\leq\kappa(\alpha,L)\,\|\Pi_{\perp}X\|_{F}.$ (25)
Proof.
Since $\Delta_{N}$ is symmetric (Lemma B.1), it admits an orthogonal eigen-decomposition $\Delta_{N}=U\Lambda U^{\top}$ with $\Lambda=\mathrm{diag}(\lambda_{0},\dots,\lambda_{L-1})$. The operator $\mathcal{S}_{\alpha}=I+\alpha\Delta_{N}$ is also symmetric with eigenvalues
$\mu_{k}=1+\alpha\lambda_{k}.$ (26)
By Lemma B.1, $\lambda_{k}\in[-4,0]$. For $0\leq\alpha<\tfrac{1}{2}$, we have
$0\leq 1-4\alpha\leq\mu_{k}\leq 1.$ (27)
Because $\mathcal{S}_{\alpha}$ is symmetric, its spectral norm equals the maximum absolute eigenvalue: $\|\mathcal{S}_{\alpha}\|_{2}=\max_{k}|\mu_{k}|\leq 1$, proving the first claim. Then for any $X$, using the standard inequality $\|MX\|_{F}\leq\|M\|_{2}\|X\|_{F}$ yields $\|\mathcal{S}_{\alpha}(X)\|_{F}\leq\|\mathcal{S}_{\alpha}\|_{2}\|X\|_{F}\leq\|X\|_{F}$.
For strict contraction off the nullspace, observe that $\mathrm{Null}(\Delta_{N})$ is spanned by the constant eigenvector (mode $k=0$), for which $\mu_{0}=1$. On $\mathrm{Null}(\Delta_{N})^{\perp}$ only modes $k\geq 1$ appear, and since $\lambda_{1}=\max_{k\geq 1}\lambda_{k}<0$ we have
$\kappa(\alpha,L)\triangleq\max_{k\geq 1}\mu_{k}=1+\alpha\lambda_{1}\in(0,1).$ (28)
Therefore $\|\Pi_{\perp}\mathcal{S}_{\alpha}(X)\|_{F}\leq\kappa(\alpha,L)\|\Pi_{\perp}X\|_{F}$. ∎
B.4 Proof of monotone Dirichlet-energy dissipation
Definition B.3 (Discrete Dirichlet energy).
For $X\in\mathbb{R}^{L\times d}$, define
$\mathcal{E}(X)\triangleq\tfrac{1}{2}\langle X,-\Delta_{N}X\rangle\;=\;\tfrac{1}{2}\sum_{j=1}^{d}X_{\cdot j}^{\top}(-\Delta_{N})X_{\cdot j}\;=\;\tfrac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}(X_{i+1,j}-X_{i,j})^{2}.$ (29)
Theorem B.4 (Monotone energy dissipation).
For $0\leq\alpha<\tfrac{1}{2}$ and all $X\in\mathbb{R}^{L\times d}$,
$\mathcal{E}(\mathcal{S}_{\alpha}(X))\leq\mathcal{E}(X),$ (30)
with strict inequality whenever $\Pi_{\perp}X\neq 0$ and $\alpha>0$.
Proof.
Because $\Delta_{N}$ acts independently across channels, it suffices to prove the claim for a single channel $x\in\mathbb{R}^{L}$ and then sum over $j$. Let $\Delta_{N}=U\Lambda U^{\top}$ and write $\hat{x}=U^{\top}x$. Then
$\mathcal{E}(x)=\tfrac{1}{2}\,x^{\top}(-\Delta_{N})x=\tfrac{1}{2}\sum_{k=0}^{L-1}(-\lambda_{k})\,\hat{x}_{k}^{2}.$ (31)
After diffusion, $\widehat{\mathcal{S}_{\alpha}x}_{k}=(1+\alpha\lambda_{k})\hat{x}_{k}$, hence
$\mathcal{E}(\mathcal{S}_{\alpha}x)=\tfrac{1}{2}\sum_{k=0}^{L-1}(-\lambda_{k})\,(1+\alpha\lambda_{k})^{2}\,\hat{x}_{k}^{2}.$ (32)
For $k=0$, $\lambda_{0}=0$ so the summand is zero. For $k\geq 1$, $\lambda_{k}>0$ and, by Eq. (27), $0\leq(1+\alpha\lambda_{k})^{2}\leq 1$, so each summand in Eq. (32) is at most the corresponding summand in Eq. (31). Summing over $k$ gives (30).
If $\alpha>0$ and $\Pi_{\perp}x\neq 0$, then there exists some $k\geq 1$ with $\hat{x}_{k}\neq 0$ and $\lambda_{k}<0$, which implies $(1+\alpha\lambda_{k})^{2}<1$ and hence strict inequality. ∎
B.5 Commutator identity and a linearized attention view
Proposition B.5 (Commutator identity).
Let $\mathcal{A}$ be any (possibly nonlinear, input-dependent) attention mapping, and let $\mathcal{S}_{\alpha}$ be the diffusion operator. Define
$Y_{\mathrm{pre}}\triangleq\mathcal{A}(\mathcal{S}_{\alpha}(X)),\qquad Y_{\mathrm{post}}\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X)).$ (33)
Then
$Y_{\mathrm{post}}-Y_{\mathrm{pre}}=[\mathcal{S}_{\alpha},\mathcal{A}]X\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}_{\alpha}(X)),$ (34)
and therefore $\|Y_{\mathrm{post}}-Y_{\mathrm{pre}}\|_{F}=\|[\mathcal{S}_{\alpha},\mathcal{A}]X\|_{F}$.
Proof.
The identity is purely algebraic by direct expansion of the definitions in Eq. (33). ∎
Linearized attention heuristic.
At a fixed input $X$, one may locally approximate attention by a mixing matrix $A(X)\in\mathbb{R}^{L\times L}$ acting along the sequence dimension (e.g., row-stochastic weights at $X$): $\mathcal{A}(X)\approx A(X)X$. Under this heuristic,
$\mathcal{S}_{\alpha},\mathcal{A}]X\approx\alpha\bigl(\Delta_{N}A(X)-A(X)\Delta_{N}\bigr)X.$ (35)
This approximation is intended as an interpretive diagnostic (not a global linear model of attention).
B.6 Lipschitz propagation bound and roughness control
Theorem B.6 (Lipschitz propagation bound).
Let $F=f_{m}\circ\cdots\circ f_{1}$ be a composition of mappings on $\mathbb{R}^{L\times d}$, and define the variant with diffusion inserted after $f_{j}$:
$F^{(j)}=f_{m}\circ\cdots\circ f_{j+1}\circ\mathcal{S}_{\alpha}\circ f_{j}\circ\cdots\circ f_{1}.$
Assume the tail maps $f_{j+1},\dots,f_{m}$ are locally Lipschitz in $\|\cdot\|_{F}$ with constants $L_{j+1},\dots,L_{m}$. For an input $X$, let $Z_{j}=(f_{j}\circ\cdots\circ f_{1})(X)$. Then
$\|F^{(j)}(X)-F(X)\|_{F}\leq\Bigl(\prod_{\ell=j+1}^{m}L_{\ell}\Bigr)\,\|\bigl(\mathcal{S}_{\alpha}-I\bigr)(Z_{j})\|_{F}.$ (36)
Moreover,
$\|\bigl(\mathcal{S}_{\alpha}-I\bigr)(Z_{j})\|_{F}=\alpha\|\Delta_{N}Z_{j}\|_{F}\leq\alpha\sqrt{2\,\mathcal{E}(Z_{j})}.$ (37)
Proof.
Define the tail composition $T\triangleq f_{m}\circ\cdots\circ f_{j+1}$. By assumption, $T$ is locally Lipschitz in $\|\cdot\|_{F}$ with constant $\prod_{\ell=j+1}^{m}L_{\ell}$. Therefore,
$\|F^{(j)}(X)-F(X)\|_{F}=\|T(\mathcal{S}_{\alpha}(Z_{j}))-T(Z_{j})\|_{F}\leq\Bigl(\prod_{\ell=j+1}^{m}L_{\ell}\Bigr)\,\|\mathcal{S}_{\alpha}(Z_{j})-Z_{j}\|_{F},$ (38)
which is Eq. (36).
From Eq. (37), we have $\mathcal{S}_{\alpha}(Z_{j})-Z_{j}=\alpha\Delta_{N}Z_{j}$. It remains to relate $\|\Delta_{N}Z_{j}\|_{F}$ to $\mathcal{E}(Z_{j})$. For a single channel vector $z\in\mathbb{R}^{L}$, Lemma B.1 implies
$\|\Delta_{N}z\|_{2}^{2}=\sum_{k}\lambda_{k}^{2}\,\hat{z}_{k}^{2}\leq 4\sum_{k}(-\lambda_{k})\,\hat{z}_{k}^{2}=8\,\mathcal{E}(z),$ (39)
because $\lambda_{k}^{2}\leq 4(-\lambda_{k})$ for all $\lambda_{k}\in[-4,0]$. Summing this bound over channels yields $\|\Delta_{N}Z_{j}\|_{F}^{2}\leq 4\,\langle Z_{j},-\Delta_{N}Z_{j}\rangle=8\,\mathcal{E}(Z_{j})$, i.e., $\|\Delta_{N}Z_{j}\|_{F}\leq\sqrt{2\,\mathcal{E}(Z_{j})}$. ∎